Given the following parameters Vin = 14 V, R1 = 24 kΩ, R2 = 12 kΩ, R3 = 6 kΩ,

R4 = 3 kΩ and a feedback resistor of 10 kΩ. Calculate the output voltage V0. (9)


1
Expert's answer
2021-08-27T02:07:36-0400

"V_o= (-\frac{R_f}{R})V_1\\"


In the first stage

"V_o= (-\frac{10*10^3}{24*10^3})*14\\
=-\frac{17.5}{3}= -5.833333V"


In the second stage

"V_o= (-\frac{10*10^3}{12*10^3})*14\\
=-\frac{35}{3}= -11.66666 V"


In the third stage

"V_o= (-\frac{10*10^3}{6*10^3})*14\\
=-\frac{70}{3}=-23.33333 V"


In the fourth stage

"V_o= (-\frac{10*10^3}{3*10^3})*14\\
=-\frac{140}{3}=-46.66666 V"